# Zero-Shot Statistical Tests for LLM-Generated Text Detection using Finite Sample Concentration Inequalities

### v3

@@ A.8 Proof of Proposition 3
where the inequality follows from Theorem 1. ∎
Appendix B Related work
@@
Supervised detection methods. Supervised detection methods learn to differentiate between human-written and LLM-generated text through labeled examples. For example, GPT2 Detector (Solaiman et al. 2019) fine-tunes RoBERTa on the output of GPT2, and ChatGPT Detector (Guo et al. 2023) fine-tunes RoBERTa on the HC3 dataset in Guo et al. 2023. Many other supervised detection methods also exist, including those that leverage neural representations (e.g., Bakhtin et al. 2019; Uchendu et al. 2020; Fagni et al. 2021) and those that leverage bag-of-words (Fagni et al. 2021). Although some supervised detection methods show high performance, it has been observed that they tend to overfit their training domains and source models (Bakhtin et al. 2019; Uchendu et al. 2020). A key concern with neural-based methods (e.g., feature-based classifiers) is their poor robustness, e.g., against ambiguous semantics (Schaaff et al. (2023)) and their limited ability to detect LLM-generated misinformation (Schuster et al. (2020)). Recent efforts have been enhancing training methodology (e.g., Kumarage et al., 2023; Tu et al., 2023). However, these methods provide heuristics without theoretical analysis to guarantee that results would hold irrespective of the specific features of their studied texts. Additionally, there are limitations for the feature-based classifiers, including challenges in training models with the hype in developing new LLMs, an increase in the variety of topics and writing styles, and legal concerns associated with training on human data, such as privacy concerns.
@@
Zero-shot detection methods. Another stream of work attempts to distinguish machine-generated from human-written texts using statistics-based methods. Most research on statistics-based methods focuses on white-box statistics, black-box statistics, and linguistics feature statistics. Among those, white-box detection methods are closely related to our work. The existing white-box detection methods primarily apply logit-based statistics or perturbed-based methods. Closer to our work are logit-based statistics. In the logit-based stream, Log-likelihood is one of the most widely used measures (Solaiman et al. 2019). Other measures include using GTLR based on rank-likelihood (Gehrmann et al. 2019), the Log-likelihood Ratio Ranking (LRR) proposed by Su et al. (2023), entropy and Kullback-Leibler (KL) divergence (Lavergne et al. 2008), and perplexity ( Vasilatos et al. 2023; Wang et al. 2023b). Mitchell et al. (2023) observes that machine-generated texts tend to lie in the local curvature of the log probability and proposes DetectGPT, which has demonstrated reliable performance but has some limitations. Du et al. analyzes the limitations of DetectGPT, highlights DetectGPT’s need for computing many perturbations, making it a computationally intensive algorithm, and proposes targeted masking strategies (rather than random masking) to improve DetectGPT’s performance. Hans et al. (2024) proposes Binoculars that use a ratio of perplexity measurement and cross-perplexity, a notion of how surprising the next token predictions of one model are to another model. Like supervised detection methods, most statistics-based detection methods provide only heuristics.
@@
Watermarking. One detection approach is to record (Krishna et al., 2024) or watermark (Kirchenbauer et al., 2023) all generated text. The method proposed in Kirchenbauer et al. (2023) takes the last generated token in the prefix and uses it to seed an RNG, which randomly places $50\%$ of the possible subsequent tokens in a green list and the remaining in a red list. During sampling, the algorithm boosts the probability of sampling a green word, resulting in a higher than $50\%$ of the final text consisting of green words. Note that watermarking requires the cooperation of the generating party/LLM owner to implement the green listing-red listing algorithm. For a survey on watermarking methods, we refer readers to Amrit and Singh (2022).
@@
Possibility of detection and robustness to attacks. Robust detection methods are being developed, and increasingly sophisticated evasion methods are being devised to circumvent these detectors, creating an ongoing contest between detection and evasion. The evasion of detection methods can be through (i) prompt engineering by general users who change the writing style of LLM-generated text or (ii) through paraphrasing attacks by adversaries. Recent research examines the ability of different detection methods against the mentioned evasion methods, including, e.g., Sadasivan et al. 2023; Krishna et al. 2024; Zhang et al. 2024. As a theoretical work in this stream, Chakraborty et al. (2023) provides (detection) possibility results for detecting machine-generated and human-generated texts. Specifically, results characterize the number of samples for the likelihood-ratio-based detector to achieve an AUROC of $\epsilon$. In binary classification, the ROC Curve is a graphical representation that illustrates the performance of a binary classification model at various thresholds, and the AUROC (Area Under the ROC Curve) quantifies model’s ability to distinguish between classes. As the decision threshold changes, AUROC shows the trade-off between the True Positive Rate (TPR) and the False Positive Rate (FPR). For a survey on possibility/impossibility of detecting AI-generated text, please refer to Ghosal et al. (2023).
@@
Authorship attribution. The literature on human-written text author attribution is extensive and, similar to the recent literature on LLM text detection and attribution, has seen developments over time, e.g., use of statistical hypothesis testing versus discriminative methods like support vector machines, neural networks, etc. In author attribution with only human authors, there is no access to the text generation "model," and the attribution is only based on the samples of texts written by different human authors. For surveys on the different methods for human author attribution, we refer readers to Juola et al. 2008; Stamatatos 2009; Koppel et al. 2009, and for a comparison of different methods, we refer readers to Grieve (2007).

### v4

@@ A.2.2 Proof of Proposition 2
where the inequality follows from Theorem 1-part (a). This concludes the proof. ∎
Appendix B Experiment details
B.1 Additional experiment results
AUROC Results. For this experiment, model samples are generated by sampling from the raw conditional distribution with temperature 1.
XSum SQuAD WritingPrompts
$\log p(x)$ 0.98* 0.84 0.99* 0.91 0.75 0.60 1.00 0.95 1.00*
Rank 0.75 0.69 0.69 0.71 0.71 0.55 0.82 0.80 0.81
LogRank 0.98* 0.87 0.99* 0.93* 0.81* 0.62 1.00 0.97* 1.00*
Entropy 0.39 0.70 0.40 0.37 0.66 0.70* 0.04 0.36 0.02
DetectGPT 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Binoculars 0.78 0.95 0.99 * 0.55 0.78 0.62 0.68 0.97* 0.99
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98* 1.00 1.00
Diff (zero-shot) 0.01 0.04 0.01 0.06 0.18 0.27 -0.02 0.03 0.00
Roberta (base) 0.98* 0.95 0.92* 0.97* 0.92 0.69* 0.97* 0.95* 0.74*
Roberta (large) 0.98* 0.98* 0.92* 0.95 0.93* 0.68 0.96 0.93 0.65
Ours 0.99 0.99 1.00 0.99 0.99 0.97 0.98 1.00 1.00
Diff (supervised) 0.01 0.01 0.08 0.02 0.06 0.28 0.01 0.05 0.26
Table 5: AUROC for detecting samples from the given models on the datasets. The best AUROC values are in bold, and the second-best values are marked with an asterisk (*). The rows Diff(zero-shot) and Diff(supervised) show our AUROC improvement over the strongest zero-shot and supervised baseline methods.
Results in Table 5 demonstrate that our method improves the average AUROC compared to the benchmarks for all datasets, with our maximum improvement obtained for the SQuAD dataset. An explanation for our higher improvement on SQuAD is the dataset’s higher entropy variance (0.128) compared to XSum (0.045) and WritingPrompts (0.074). Since our method relies on the difference between log-perplexity and the average cross-entropy, it performs more effectively when entropy’s variance is high, while aligning with likelihood-based methods when the entropy is less variable. Another observation is that our method significantly outperforms existing likelihood-based approaches in detecting Qwen-generated text for SQuAD. An explanation is that QWEN 32B is a larger model with more human-like syntax and fewer low-entropy patterns.
Convergence results. We conduct our numerical analysis for convergence on the following pre-trained language models: GPT-2 small, GPT-2 medium, GPT-2 large, GPT-2 XL, and GPT-Neo. Through experiments with different generative and evaluator models, we examine whether the log-perplexity of a short portion of text converges to the average cross-entropy. Our experiments measure these values across generated text and analyze their performance over different configurations. Our setup includes generating tokens with pre-trained models and recording each token’s selection probability and calculated metrics.
Same generative and evaluator model. In the first set of experiments, we employ GPT-2 to generate a series of $100$ tokens, beginning with the fixed prompt "Jack". We use the model’s conditional probability distribution for each token generation step to sample the next token. Note that for the white-box model of GPT-2, probability distributions are accessible. We calculate each generated token’s empirical entropy and log-perplexity and repeat this process for comparisons. We use Softmax-normalized probabilities to select the next token and store the generated token and its probability distribution. For each sub-string of length $N$ starting from the first token in the generated sequence, we compute the log-perplexity $l_{A}(\mathbf{Y}_{N})$, and the empirical entropy $h_{N}(A,A)(\mathbf{Y}_{N})$. The results are shown in Figures (1(a)-1(d)). We consistently observe that the numerical results confirm Theorem 1- part (a) that the log-perplexity converges to the average entropy when the generative and evaluator models are the same.
(a) GPT-2 small
(b) GPT-2 medium
(c) GPT-2 large
(d) GPT-2 XL
Figure 1: Generated and evaluated by the same model
Different generative and evaluator models. To extend our numerical analysis to the case with different generative and evaluator models, we generate a string using the following generative models: GPT-2 medium, GPT-2 large, and GPT-2 XL. Then, we calculate the log-perplexity of these strings using the evaluator model GPT2-small. We calculate the cross-entropy of the strings under each generative model and the evaluator model (GPT2- small). The results are shown in Figures (2(a)-2(c)). Results in these figures confirm Theorem 1. In particular, we observe that when the evaluator and generative models are different, the log-perplexity of the string converges to the average cross-entropy of the string under generative and evaluator models.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XL
Figure 2: Generated by GPT-2 small and evaluated by different models
B.2 Dataset details
We use the codes in the official GitHub repository for DetectGPT [42], which is publicly available under the MIT license in https://github.com/eric-mitchell/detect-gpt for our white-box experiments. The datasets for this set of experiments are XSum, SQuAD, and Reddit writingPrompts that we briefly describe below. We use the codes in the official GitHub repository for the RAID dataset [11] for our black-box experiments. The RAID dataset uses a variety of datasets from 8 domains, briefly described below. For details about the RAID dataset, please refer to Appendix E.1 in [11].
XSum.
The Extreme Summarization (XSum) dataset comprises 226,711 BBC news articles paired with single-sentence summaries for each article. Each entry includes a document (the full news article), a summary (a concise, one-sentence abstraction), and an id (the BBC article identifier). The dataset is designed to support the task of extreme abstractive summarization, where models must generate highly condensed summaries that capture the essence of the source text. The XSum dataset is publicly available under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license and can be accessed at https://huggingface.co/datasets/EdinburghNLP/xsum. In our experiments, we utilize the first 30 tokens of the document column as input for creating LLM-generated text.
SQuAD.
The Stanford Question Answering Dataset (SQuAD) is a benchmark for machine reading comprehension, consisting of over 100,000 question-answer pairs derived from 536 Wikipedia articles. Each example includes a context paragraph, a question, and one or more answers, where each answer is a span of text from the corresponding context. SQuAD is publicly available under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) and can be accessed at https://huggingface.co/datasets/rajpurkar/squad. In our experiments, we utilize the first 30 tokens of the context column as input for creating LLM-generated text.
Reddit WritingPrompts.
The Reddit WritingPrompts dataset comprises over 300,000 prompt-story pairs collected from the r/WritingPrompts subreddit. Each entry includes a prompt (a creative writing prompt) and a corresponding story (a user-generated narrative response). The dataset is designed to support research in open-ended story generation and narrative modeling. The dataset is publicly available under the MIT License and can be accessed at https://huggingface.co/datasets/euclaise/writingprompts. In our experiments, we utilize the first 30 tokens of the story field as input for creating LLM-generated text.
RAID.
The Robust AI Detection (RAID) dataset is a large-scale benchmark constructed to evaluate the reliability of LLM-generated text detectors across a wide range of textual domains. It comprises over 6 million generations created by 11 language models under 4 decoding strategies and 11 adversarial attacks. RAID includes 8 carefully selected text domains to capture a broad spectrum of natural language styles, structures, and levels of creativity. The domains are
Abstracts. Sourced from arXiv.org, this domain consists of paper abstracts across multiple disciplines. Texts are technical, information-dense, and syntactically formal, reflecting real-world scientific writing in academic publishing.
Books. Drawn from the BookCorpus dataset, this domain features long-form fictional passages including dialogue and narration. It captures informal, imaginative writing with strong character and plot structure diversity.
News. Extracted from the CNN/DailyMail summarization corpus, the news domain includes professionally written news stories. These texts are structured with leads and summaries, and often exhibit a neutral tone, making them ideal for assessing generation consistency in journalistic formats.
Poetry. Taken from Project Gutenberg, the poetry domain includes English poems from public domain authors. The texts exhibit high stylistic variability, creative syntax, line breaks, and metaphorical expressions, which are characteristics that typically challenge detection.
Recipes. Based on the RecipeNLG dataset, this domain contains step-by-step cooking instructions and ingredient lists. It features a highly structured and instructional tone with imperative phrasing and domain-specific vocabulary.
Reddit. Sourced from the WritingPrompts subreddit, these texts are user-generated responses to open-ended prompts. The domain features creative fiction in a casual tone, reflecting informal narrative style and internet native conventions.
Reviews. Pulled from Amazon product reviews, this domain contains subjective, opinionated writing with informal grammar and varying tone. It reflects real-world user-generated content that combines evaluation, anecdote, and justification.
Wikipedia. This domain includes encyclopedic entries from Wikipedia, selected for their balanced, objective tone, structured format, and high factual density. Texts are varied in topic and are typically well-edited and linguistically standardized.
To avoid data contamination, all human-written documents were sourced from publicly available corpora with timestamps preceding 2022. For each document, the dataset provides one generation per model, decoding strategy, and adversarial attack. The dataset is publicly available at https://huggingface.co/datasets/liamdugan/raid under the MIT License.
B.3 The initial thirty human tokens’ impacts
Recall that for our whitebox experiments in Section 5.2, we randomly sample 500 texts from each dataset, and prompt the first 30 tokens of each text to the LLM to create an LLM-generated version. Note that the first 30 tokens in both samples by human and LLM are the same. Thus, these 30 tokens result in the same average entropy and log perplexity under both null and alternative. Denote the random variable by $C=\sum_{i=1}^{30}X_{n}$. Also, since a human generated the first 30 tokens the random variable $C$ is not zero mean. Hence, note that for both null and alternative $\sum_{n=1}^{300}X_{n}$ is biased by $C$ which is not a zero mean random variable. So, under this setting, it is not appropriate to check $|\sum_{n=1}^{300}X_{n}|$ is small to determine null. On the other hand, if we use the one-sided version of our test that rejects the null hypothesis when $\frac{1}{N}\sum_{n=1}^{N}X_{n}^{A}=\frac{1}{N}\sum_{n=1}^{N}Z_{n}^{A}-h_{N}(A,A)(\mathbf{Y}_{N})\leq t.$ Since, the $\sum_{i=1}^{300}X_{i}$ under null and alternative is shifted by a constant $C$. The performance of the test will not be affected. Hence, we use the one-sided version of our one-sided test that removes the impact of the differences between cross-entropy and log perplexity caused by the initial 30 tokens.
B.4 Decoding strategies & Adversarial attacks
B.4.1 Decoding strategies
A decoding strategy determines how tokens are selected from the LLM’s (conditional) probability distribution in the text generation process. Greedy decoding selects the most likely (top-ranked) token from the (conditional) probability distribution. Previous research has shown that sampling from the probability distribution rather than selecting the top-ranked token (as done in greedy decoding) can make the LLM-generated text less detectable [28]. Based on these findings, [11] generates two outputs per prompt: one with greedy decoding and the other with random sampling.
The repetition penalty, as defined in [32], reduces the likelihood of generating previously used tokens through down-weighting the probability of choosing the same word in the following tokens using a parameter $\theta$. This penalty examines whether the less-repetitive output (that has undergone repetition penalty) can reduce the detectability of LLM-generated text. Formally, Given a temperature $T>0$ and a set of scores $s_{i}\in\mathbb{R}^{d}$ for each token $i$ in a vocabulary, the probability $p_{i}$ of predicting the $i^{\text{th}}$ token is
$p_{i}=\frac{\exp(s_{i}/T)}{\sum_{j}\exp(s_{j}/T)}.$
The repetition penalty defined by [32] modifies the distribution as
$\tilde{p}_{i}=\frac{\exp\left(\frac{s_{i}}{T\cdot\mathbb{I}(i\in g)}\right)}{\sum_{j}\exp\left(\frac{s_{j}}{T\cdot\mathbb{I}(j\in g)}\right)},$
where $g$ is the list of previously generated tokens, and the indicator function $\mathbb{I}(c)$ is
$\mathbb{I}(c)=\begin{cases}\theta&\text{if }c\text{ is true},\\ 1&\text{otherwise}.\end{cases}$
[11] uses [32]’s method and generates two additional outputs: with and without repetition penalty. Following [32], we use $\theta$ = 1.2 for the experiments.
B.4.2 Adversarial attacks
Adversarial attacks in LLM detection are strategically crafted paraphrases or perturbations applied to LLM-generated text that aim to bypass detection by reducing detectable differences between LLM-generated and human-written text. [11] produces different adversarially attacked versions of texts for each LLM-generated text. Below are brief descriptions of the attacks and sources from which [11] acquired them.
Insert Paragraphs [5]
[…]
Article Deletion [40, 21]
first searches through the text and captures all articles "the," "a," and "an" using POS tagging. Then, using random sampling, selects a fixed $\theta$ percent of the articles and applies the article deletion to them.
Alternative Spelling [41]
constructs a mapping between British and American spelling of words using an American to British English dictionary (https://github.com/hyperreality/American-British-English-Translator). Then, searches in the LLM-generated text to identify all the words (that also exist) in the dictionary, randomly samples a fixed $\theta$ percent of the words (with two spellings), and replaces them with the alternative spelling.
Misspelling [40, 14]
uses Wikipedia’s commonly misspelled English words dictionary (https://en.wikipedia.org/wiki/Commonly_misspelled_English_words), and finds the LLM-generated words that also exist in the misspelling dictionary.
Upper Lower Swap [14]
randomly selects a fixed $\theta$ percent of the words in the LLM-generated text, and changes the first letter to uppercase if originally lowercase, and vice versa.
Number Swap [5]
first searches through the text, and captures all numerical digits in the LLM-generated text. Then, randomly selects fixed $\theta$ percent of the digits and replace them with a randomly selected digit from 0 to 9.
Whitespace Addition [6, 14]
captures all the spaces between tokens, randomly selects a fixed $\theta$ percent of the spaces, and inserts one additional space to each selected space. Note that since the sampling is performed with replacement, the same inter-token space can be chosen multiple times, resulting in a possible insertion of more than one extra space between tokens.
Paraphrasing [35, 52]
transforms the source LLM-generated text into an alternative text with the same semantic content (i.e., the conveyed meaning is equivalent, but the content is statistically and stylistically different). [11] uses the paraphrasing method in [35]’s DIPPER-11B using their HuggingFace repository. DIPPER-11B is a paraphrasing model built on the T5-XXL architecture. The process includes two control parameters: one for lexical diversity (a measure of the variety of the words used in the text) and another for reordering, which adjusts syntactic transformation. DIPPER-11B encodes the LLM-generated text as input and then decodes a paraphrase as output. [11] uses the default setting from [35] to apply the paraphrasing attack on all texts in the dataset.
Synonym Swap [49]
does the following process (as done in [11]): first, they mask all the tokens, replacing each with a mask-fill from the top-twenty candidates in BERT [9]. Next, they assign the Part-of-Speech (POS) tag, which is a label assigned to each word in a sentence to indicate its grammatical category (such as noun, verb, adjective, or adverb), for each of the mask-fill candidates and remove all those candidates whose POS tag does not match the POS tag for the token in the input text. After creating the final list of valid candidate swaps, they select a fixed $\theta$ percent of the texts for performing the Synonym Swap adversarial attack.
Homoglyph [70, 14]
swaps some English characters with visually similar yet technically different characters from the Cyrillic scripts in [70]. Specifically, [11] uses homoglyphs for the following ASCII characters: a, A, B, e, E, c, p, K, O, P, M, H, T, X, C, y, o, x, I, i, N, and Z. They identify all these characters in the LLM-generated text, and then replace all of them (i.e., $\theta$=100%) with the homoglyph counterparts. If a character has more than one homoglyph, they randomly select between them.
Zero-width Space [21]
the Unicode zero-width space character U+200B is an invisible character that occupies space in the underlying text encoding without being visible to human readers. [11] inserts a zero-width space before and after every visible character in the perturbed text.
Note that [11]’s dataset for the Homoglyph and Zero-width space attacks is incomplete, not allowing us to provide results for them. Nevertheless, these two attacks are easy to remove with a simple pre-processing that we do in our codes (details in Appendix B.7): First, remove or normalize the invisible characters like zero-width spaces. Then map suspicious non-ASCII characters to standard ASCII or a canonical form.
To find the $\theta$ that [11] uses for each adversarial attack, please refer to Table 11 in their paper.
B.5 Detector details
RAID [11] evaluates 12 detection methods from neural classifiers, metric-based (zero-shot) detectors, and commercial detectors. All detectors were evaluated using their official implementations or web interfaces without fine-tuning on RAID data. Below is a brief description of each detector and how it was run in [11]. To check how [11] finds the thresholds to achieve FPR of 5%, please refer to Appendix F-2 in [11].
Neural classifier-based
RoBERTa (GPT2): fine-tuned to distinguish between human and GPT-2-generated text. It was evaluated using the released Hugging Face checkpoint without modification. RoBERTa-B (GPT2) is a smaller RoBERTa-base version trained on the same dataset. It is computationally lighter, but typically less accurate than the large model.
RoBERTa-B (ChatGPT) (22) fine-tuned to detect generations from ChatGPT. [11] downloads and queries the detector via HuggingFace datasets with the unique identifier Hello-SimpleAI/chatgpt-detector-roberta.
RADAR (25) a robust RoBERTa-based detector adversarially trained to resist paraphrasing and rewriting attacks. Used as released without retraining. [11] downloads and queries the detector via HuggingFace datasets with the unique identifier TrustSafeAI/RADAR-Vicuna-7B.
Metric-based (Zero-shot)
GLTR (16) evaluates the likelihood of text according to an (evaluator) LLM, bins tokens based on their likelihoods, and uses the bins as signals for detection. [11] uses the default settings from the GLTR repository (https://github.com/HendrikStrobelt/detecting-fake-text), which is setting the (evaluator) LLM to GPT2 small and the cutoff threshold at rank=10.
Fast DetectGPT (4) DetectGPT evaluates log-probability curvature under paraphrased perturbations. Low curvature regions suggest LLM generations. It was implemented using DetectGPT’s official repository with standard settings. A more efficient variant of DetectGPT that avoids expensive perturbation by approximating curvature metrics is Fast DetectGPT [4]. It was implemented using the default repository configuration, GPTNeo-2.7B as the scoring model, and GPT-J-7 B as the reference model.
Binoculars (24) uses perplexity divided by cross-perplexity for detection. RAID repository uses the code from the Binocular GitHub repository and calculates perplexity using the default models from Binocular’s repository, which are Falcon 7B and Falcon 7B Instruct.
LLMDet (71) uses proxy-perplexity, an approximation of perplexity calculated by repeatedly sampling n-grams from models rather than running the models, of the input text from 10 different small language models. None of the evaluator models used in the detector was used for text generation.
Commercial
GPTZero
([59]) relies on two metrics: perplexity, which measures how predictable the text is under an evaluator LLM, and burstiness, which captures the variation in perplexity across sentences. Human-written text typically exhibits higher burstiness, while LLM-generated text tends to be more predictable [60]. [11] queried the detector using the v2 API (https://api.gptzero.me/v2/predict/text) and threshold on the completely_generated_prob field.
Originality
[11] queried this detector through the v1 API (https://api.originality.ai/api/v1/scan/ai) and threshold on the score field in the output JSON. They queried the multilingual "version 3” of the detector for the Czech and German news domains.
Winston
[11] queried this model through the v1 API (https://api.gowinston.ai/functions/v1/predict), set the input language to English, unless the text is in German, where they set the language to German. They threshold on the score field in the output JSON.
ZeroGPT
[11] queried the API (https://api.zerogpt.com/api/detect/detectText), and used the isHuman field as the classifier output.
B.6 Hardware
All experiments were conducted using a RunPod cloud instance configured with a B200 GPU, 28 vCPUs, and 283 GB of RAM. The setup provided 150 GB of disk and pod volume and achieved bandwidth speeds of up to 20,801 Mbps in the EU-RO-1 region. Experiments were executed in an on-demand secure cloud environment available at https://www.runpod.io. Our complete set of experiments required a total of approximately 12 hours of runtime. Because our method is nonparametric and does not require training, the runtime scales linearly with the number of input texts in the dataset. Figure 3 demonstrates the RunPod cloud configuration.
Figure 3: RunPod cloud configuration used for all experiments.
B.7 Code instructions
Contents
We include our self-contained codes as part of our supplementary materials to enable replication of all experiments presented in this work. The repository comprises three Jupyter notebooks: Whiteboxattribution.ipynb for identifying the generating model among two candidate sets (codes for the experiments in Section 5.1), Whiteboxdetection.ipynb for statistical detection tests in a white-box setting (codes for the experiments in Section 5.2), and Blackbox_adversarial.ipynb for evaluating black-box detection and robustness against adversarial attacks (codes for the experiments in Section 5.3). All experiments were run on RunPod (B200) and used Hugging Face models and datasets where applicable.
How to run
Each notebook is fully self-contained and does not require external data downloads or model training, facilitating straightforward execution on personal devices. Each notebook includes
  • Installation and environment setup (in the first cells)
  • Data loading and preparation
  • Full experiment code with results
No setup beyond running the notebook cells is required since all dependencies are installed inline.
Datasets
Most datasets used in the experiments are downloaded programmatically from Hugging Face.
However, if you want to run the WritingPrompts experiments, you will need to download the WritingPrompts dataset manually from https://github.com/facebookresearch/WritingPrompts. Then, you need to place the downloaded data in the following path: data/writingPrompts/.
Portions of this code are adapted from DetectGPT, available at https://github.com/eric-mitchell/detect-gpt/tree/main. We thank the authors of DetectGPT for open-sourcing their codebase.
Appendix C Literature review
Neural classifier-based detection methods.
@@ Appendix C Literature review
Neural classifier-based methods learn to differentiate between human-written and LLM-generated text through labeled examples of human-written and LLM-generated text in the training set. For example, GPT2 Detector (Solaiman et al. [56]) fine-tunes RoBERTa on the output of GPT2, and ChatGPT Detector (Guo et al. [22]) fine-tunes RoBERTa on the HC3 dataset in Guo et al. [22]. Many other supervised detection methods also exist, including those that leverage neural representations (e.g., Bakhtin et al. [3], Uchendu et al. [63]) and those that leverage bag-of-words (Fagni et al. [12]). Although some supervised detection methods show high performance, it has been observed that they tend to overfit their training domains and source models (Bakhtin et al. [3], Uchendu et al. [63]). A key concern with neural-based methods (e.g., feature-based classifiers) is their poor robustness, e.g., against ambiguous semantics (Schaaff et al. [53]), their limited ability to detect LLM-generated misinformation (Schuster et al. [54]), and their limited generalizability. Liang et al. [41] notes that since (neural-based) detectors are often evaluated on simple datasets, their performance on out-of-domain samples is typically abysmal. For example, they state that TOEFL essays written by non-native (human) English speakers were mistakenly marked as LLM-generated in 48-76% of detection attempts using neural-based detectors. In summary, there are limitations for the neural classifier-based methods, including challenges in training models with the hype in the development of new LLMs, an increase in the variety of topics and writing styles, and legal concerns associated with training on human data, such as privacy concerns. Recent efforts have been enhancing training methodology (e.g., Kumarage et al. [37], Tu et al. [62]). However, these methods provide heuristics without theoretical analysis to guarantee that results would hold irrespective of the specific features of their studied texts.
Metric-based/ Zero-shot detection methods.
@@
Another stream of work attempts to distinguish machine-generated from human-written texts using statistics-based methods. Most research on statistics-based methods focuses on white-box statistics, black-box statistics, and linguistics feature statistics. Among those, white-box detection methods are closely related to our work. The existing white-box detection methods primarily apply logit-based statistics or perturbed-based methods. Closer to our work are logit-based statistics. In the logit-based stream, Log-likelihood is one of the most widely used measures, e.g., as in Solaiman et al. [56]. Other measures include using GTLR based on rank-likelihood (Gehrmann et al. [16]), the Log-likelihood Ratio Ranking (LRR) proposed by Su et al. [58], entropy and Kullback-Leibler (KL) divergence (Lavergne et al. [38]), and perplexity (Vasilatos et al. [65], Wang et al. [68]). In a more recent paper, Mitchell et al. [42] observes that machine-generated texts tend to lie in the local curvature of the log probability and proposes DetectGPT, which has demonstrated reliable performance but has some limitations. Du et al. [10] analyzes the limitations of DetectGPT, highlights DetectGPT’s need for computing many perturbations, making it a computationally intensive algorithm, and proposes targeted masking strategies (rather than random masking) to improve DetectGPT’s performance. To reduce the DetectGPT’s perturbation computation cost, Bao et al. [4] introduces FastDetectGPT, which substitutes DetectGPT’s perturbation step with Conditional Probability Curvature (CPC) that accelerates the detection process by a factor of 340. Most methods focus on a single statistical property. Hans et al. [24], however, introduces a normalized measure of perplexity by dividing log-perplexity by average cross-entropy. The rationale is that LLM text is predictable to the LLM (low perplexity), whereas human text is more surprising (higher perplexity). Yet, consideration of high perplexity as a human authorship signal is unreliable due to prompt dependency [^4] This is illustrated by the ”Capybara Problem” in Hans et al. [24]. Given the prompt ”Can you write about a Capybara astrophysicist?”, an LLM may generate ”The Capybara studied dark matter in Andromeda.” Without the prompt, ”capybara” and ”astrophysicist” seem unexpectedly high in perplexity, falsely suggesting human authorship.. Similar to supervised detection methods, most statistics-based detection methods provide only heuristics.
Watermarking.
@@
Another detection approach is to record (Krishna et al. [36]) or watermark (Kirchenbauer et al. [33]) all generated text. The method proposed in Kirchenbauer et al. [33] takes the last generated token in the prefix and uses it to seed an RNG, which randomly places $50\%$ of the possible subsequent tokens in a green list and the remaining in a red list. During sampling, the algorithm boosts the probability of sampling a green word, resulting in a higher than $50\%$ of the final text consisting of green words. Note that watermarking requires the cooperation of the AI company/LLM owner to implement the green listing-red listing algorithm. For a survey on watermarking methods, we refer readers to Amrit and Singh [1].
Robustness to adversarial attacks.
Growing literature examines the susceptibility of LLM detectors to adversarial attacks. Krishna et al. [35] introduces DIPPER-11B, showing that controlled paraphrasing can significantly reduce detection accuracy for different methods, e.g., DetectGPT and GPTZero. Gagiano et al. [14] benchmarks multiple detectors under a wide range of black-box attacks, including backtranslation, synonym swaps, and whitespace addition, and reports consistent detection failure across commercial and open-source detectors. Ippolito et al. [28] demonstrates that simple perturbations, such as changing punctuation or adding unrelated phrases, helps bypass classifiers like GPT-2 detectors and GLTR. Sadasivan et al. [52] proposes multiple realistic adversarial perturbations, including paraphrasing and word deletion and finds significant performance degradation of detectors under such adversarial attacks.
Possibility of detection.
@@
As a theoretical work on the possibility of detection, Chakraborty et al. [8] provides (detection) possibility results for detecting machine-generated and human-generated texts. Specifically, results characterize the number of samples for the likelihood-ratio-based detector to achieve an AUROC of $\epsilon$  [^5] Recall that in a binary binary classification, the ROC Curve is a graphical representation that illustrates the performance of a binary classification model at various thresholds, and the AUROC (Area Under the ROC Curve) quantifies the model’s ability to distinguish between classes. As the decision threshold changes, AUROC shows the trade-off between the True Positive Rate (TPR) and the False Positive Rate (FPR).. For a survey on the possibility/impossibility of detecting AI-generated text, please refer to Ghosal et al. [17].
Authorship attribution.
@@
The literature on human-written text author attribution is extensive and, similar to the recent literature on LLM text detection and attribution, has seen developments over time, e.g., use of statistical hypothesis testing versus discriminative methods like support vector machines, neural networks, etc. In author attribution with only human authors, there is no access to the text generation "model," and the attribution is only based on the samples of texts written by different human authors. For surveys on the different methods for human author attribution, we refer readers to Juola et al. 31, Stamatatos 57 and Koppel et al. [34]. For a comparison of different methods, we refer readers to Grieve [20].